### v1

@@ Abstract
Nordic skiing provides fascinating opportunities for mathematical modelling studies that exploit methods and insights from physics, applied mathematics, data analysis, scientific computing and sports science. A typical ski course winds over varied terrain with frequent changes in elevation and direction, and so its geometry is naturally described by a three-dimensional space curve. The skier travels along a course under the influence of various forces, and their dynamics can be described using a nonlinear system of ordinary differential equations (ODEs) that are derived from Newton’s laws of motion. We develop an algorithm for solving the governing equations that combines Hermite spline interpolation, numerical quadrature and a high-order ODE solver. Numerical simulations are compared with measurements of skiers on actual courses to demonstrate the effectiveness of the model.
@@ 1 Background
  • We replace the piecewise linear course geometry from MSH [19] with a smoother Hermite cubic spline interpolant that more accurately captures the shape of a real ski course.
@@
  • The governing equations are solved using a state-of-the art numerical ODE solver (Matlab’s ode15s) that incorporates adaptive time stepping, error control and event detection.
@@
  • We incorporate a realistic 3D course that includes track curvature, which allows us to simulate different braking techniques used by athletes to safely negotiate tightly-curved downhill sections [3]. This is important in race settings where athletes strive for any possible advantage over their competitors.
@@
Exercises, Matlab Code and Data Files
@@
All data files and Matlab codes used to generate the numerical results are also provided with the Supplementary Materials.
@@ 2.1 Parameterizing a Ski Course
The path that a skier follows while propelling themselves along a real ski course is a 3D curve such as that depicted in Fig. 1a. This particular plot is generated from GPS data [6] measured on the 4.2 km “Ole” course in Toblach, Italy. This course has been officially certified for competition purposes by FIS (Fédération Internationale de Ski) and has been the location of many high-level races including the 2024 FIS World Cup Tour de Ski. The raw GPS data consists of 58 measurements of latitude $\lambda$, longitude $\psi$, and elevation $z$, which must first be converted to Cartesian coordinates on a flat patch of the Earth’s surface. The curve in Fig. 1a is plotted in terms of $(x,y,z)$, where $z$ is elevation, and $x,y$ lie in the horizontal plane [^4] When latitude $\lambda$ and longitude $\psi$ are measured in degrees, the north-south distance (in m) is $y\approx 111133\psi$ whereas the east-west distance is $x\approx 111413\lambda\cos\left(\pi\lambda/180\right)$. The latter contains an extra cosine adjustment factor because the distance between lines of longitude varies with $\lambda$. These two formulas are actually the leading order terms in an approximation that accounts for the oblate spheroidal shape of the Earth, as explained in the Wikipedia article “Geographic coordinate system.”. For convenience, we also translate coordinates so that the course starts at the origin (highlighted in the plot by a green square). A second view of the course is obtained in Fig. 1b by “unwinding” or straightening this 3D curve and displaying it as a 2D plot of elevation versus a new distance variable $\xi$, which represents the arc length measured along the projected path in the $x,y$-plane (the dotted magenta curve in Fig. 1a). This 2D elevation plot of $(\xi,z)$ is the official method for reporting course elevation profiles in competition documents (private communication, A. Maddox, June 9, 2024).
@@ 2.2 Interpolating with Cubic Hermite Splines
With a bit of foresight, it should be clear that any model of Nordic skiing will require quantities such as slope (or inclination angle) and curvature at the skier’s position $\vec{r}(\alpha)$, parameterized by $\alpha\in[0,1]$. Slope and curvature depend on the first two derivatives of position, $\vec{r}_{\alpha}$ and $\vec{r}_{\alpha\alpha}$, so that any interpolant we use must be twice differentiable. This suggests using a spline interpolant, which is a piecewise-defined function that interpolates each successive pair of data points with a polynomial. For a 2D course a single interpolant $z=Z(\alpha)$ will suffice, but in 3D we will need to construct three spline interpolants, one for each of the coordinate functions: $x=X(\alpha)$, $y=Y(\alpha)$ and $z=Z(\alpha)$.
@@ 2.3 Computing Distance and Inclination Angle
[…]
@@
[…]
When implementing these distance calculations in Matlab, the integrals in (1) and (2) must be computed numerically. To obtain accurate approximations for $s$ (and also $\xi$ for 3D courses), the coordinate splines ($\Xi,Z$ in 2D, $X,Y,Z$ in 3D) are first evaluated on a fine grid of $\alpha$ points. For the 4.2 km Ole course, we use 840 points which yields an average resolution of 5 m for each $\alpha$-interval. The integrals are then approximated using Simpson’s rule, which is fourth-order accurate and has a particularly simple implementation when the $\alpha$ points are equally spaced (see the Matlab code cumsimpson.m [^7] Conveniently, the spline data structure used in Matlab permits spline derivatives $X_{\alpha}$, $Y_{\alpha}$, $Z_{\alpha}$ appearing in the arc length integrands to be computed exactly using the function fnder.). Simpson’s rule is preferable to Matlab’s built-in trapezoidal rule function trapz because of its much higher accuracy. Once the corresponding fine-grid values of $\xi$ are computed, we can then re-parameterize the curve in terms of $\xi$ by computing a new spline for $Z(\xi)$ (plus $X(\xi)$ and $Y(\xi)$ in 3D). Indeed, the smooth elevation curve plotted in Fig. 1b was generated by first approximating the integral (1) with Simpson’s rule, and then fitting a Hermite spline to the resulting $\xi,z$ data.
@@
Because we have already built the spline $Z(\xi)$ for elevation, the derivative $Z_{\xi}$ can be computed easily using Matlab’s fnder. Then $\theta$ can be evaluated at spline points and used to construct another spline $\Theta(\xi)$ for the inclination angle. This procedure is implemented in the Matlab code setup2d.m, which constructs the Hermite splines based on 2D elevation data from a CSV file. More details are provided in Section 5 regarding the corresponding code setup3d.m for constructing a 3D parameteric spline representation from GPS data.
@@
The model devloped by MSH [19] parameterizes a ski course as $z(s)$ using arc length $s$, and replaces slope $\frac{dz}{d\xi}$ in Eq. (3) with $\frac{dz}{ds}\big{/}\sqrt{1-(\frac{dz}{ds})^{2}}$. Show that these two expressions for slope are equivalent.
@@ 2.4 Homologation and Data Errors
All ski courses that run high-level Nordic races must adhere to guidelines set out in the FIS Cross-Country Homologation Manual [8]. According to the dictionary, “homologation” is the granting of approval by some official authority. In the context of skiing, homologation refers to the FIS-certified process that any race course must go through before an official competition can be held. The manual sets out constraints on course terrain and stadium layout, including limits on the number, length, and slope of hills that are allowed [^8] An easily-digestible summary of the homologation rules can be found on the Wikipedia page “Cross-country skiing trail”.. These specifications are implemented in the form of an Excel spreadsheet that FIS officials use to perform the necessary calculations [8].
@@
    To ensure athlete safety, a maximum limit on curvature of downhill sections is imposed in terms of centripetal acceleration, $a_{c}=v^{2}/R$, where $v$ is the skier speed and $R$ is the radius of curvature. A table in [8, p. 26] identifies various limits on $a_{c}$ ranging from 2–25 m/s2 depending on race format.
@@ 3.1 ODE System for Skier Dynamics
    where $m$ is the skier mass (in kg) and $P(v)$ represents the power (in watts or kg m2/s3) that is exerted by the muscles to generate forward motion. The “$\frac{1}{v}$” dependence implies that the skier works hardest to accelerate at low speeds, whereas the power demands on their muscles ease off at higher speeds [^10] Our earlier assumption that $v>0$ ensures that this propulsion force is well-defined. However, we could handle a zero speed by replacing the factor $v$ in the denominator of (4) with $\max(v,\varepsilon)$ for some small cut-off parameter $\varepsilon>0$.. This function was proposed by MSH [19] based on a fit to measurements of Nordic ski racers and the specific form for $P(v)$ will be described later in Section 3.3.
@@
  • Aerodynamic drag force $F_{d}(v)=-\beta v^{2}$: which we assume depends quadratically on the speed relative to still air [^11] If there is a significant headwind/tailwind, then $v$ should be replaced with a wind-corrected speed (see Exercise 15).. The parameter $\beta=\frac{\rho C_{d}A}{2m}$ (units of m-1) is proportional to air density $\rho$ (in kg/m3), skier cross-sectional area $A$ (m2), and a dimensionless drag coefficient $C_{d}$. As their speed varies throughout a race, trained skiers will adjust their technique and body orientation so that the area $A$ varies with the skier’s speed. So in practice, we will consider the product $C_{d}A(v)$ to be a given function of speed (discussed in more detail in Section 3.3).
@@
The inclination angle appearing in the right hand side of both ODEs can be obtained by simply evaluating the spline $\theta=\Theta(\xi)$ constructed in Section 2.3.
@@
Use dimensional analysis (by applying the Buckingham Pi Theorem) to derive a general form for the propulsion force in Eq. (4), assuming only that $F_{p}$ depends on the quantities ${P}_{\text{\emph{max}}}$, $m$ and $v$.
@@
    Find an explicit formula for speed $v(t)$ when the course profile is $z(\xi)=b\xi$ for some constant slope $b$. Assume that the skier applies a constant propulsion force, $F_{p}(v)\equiv F_{po}$, which implies that the MSH power function is linear in the speed, $P(v)\propto v$. In constrast with Eq. (4), this force has no “$\frac{1}{v}$” singularity, so you may also assume that the skier starts from rest with $v(0)=0$.
@@
    Plot the speed $v(t)$ for an uphill climb with a gradient of $+$10% ($b=0.1$) and assume parameter values $\beta=0.0045$, $g=9.8$, $\mu=0.037$, and $F_{po}=1.4$ from the MSH baseline test (see Table 1 and Fig. 5). Observe that the skier’s speed initially increases but then gradually flattens out over time and approaches a constant terminal value. Estimate this terminal speed and the time required to reach it.
@@
    Repeat part (b) on a constant downhill gradient of --10%, assuming the skier rests in a tuck position with $\beta=0.0025$ and $F_{po}=0$. Compare your results for the uphill and downhill slopes.
@@ 3.3 Fitting Parameters to Skier Measurements
The final remaining detail is the locomotive power function $P(v)$ introduced in Eq. (4). MSH used extensive measurements of ski racers to propose three different functional forms for $P(v)$, which they compared using numerical simulations. We will adopt their “Model 1” that they used in their baseline simulations
$P(v)={P}_{\text{\emph{max}}}\exp\big{(}\text{--}\,(v/b)^{n}\big{)},$ (10)
which is designed to mimic an athlete who generates the highest power when skiing uphill (at lower speeds) and the lowest on downhills (at high speeds). This reflects the same adjustments in technique for the drag coefficient, where the skier has to use their entire body to work hardest on steep uphill sections and maintain their speed, whereas on steep downhills they shift into a tuck position to conserve energy and minimize drag. An intermediate level of power is expended on undulating terrain, where a relatively high speed can be maintained using less energy-intensive techniques. The parameters in this power equation were estimated by fitting with experimental data to obtain $n=4$, $b=10$, and ${P}_{\text{\emph{max}}}=275$.
@@
The system of nonlinear ODEs (5)–(6), or its augmented formulation (5)–(8), can be solved easily and accurately using Matlab’s stiff, adaptive ODE solver ode15s. This has been implemented in the code skirun2d.m with the variable neqs set to either 2 or 4 (the number of ODEs) to select the desired formulation. In the following sections, we apply the code to solving three test problems:
@@
    The MSH baseline simulation [19] on a 4.2 km course using the reduced (2-ODE) model. This provides a convenient test to validate our Matlab code and investigate any differences that arise due to our choice of ODE solver and spline interpolant (linear or cubmic Hermite).
@@
A major advantage of ode15s is that it adaptively adjusts the time step and the order of the numerical method to satisfy user-specified error tolerances (we use AbsTol = RelTol = 1e-8). This ensures that especially in the more difficult sections of the course (where curvature is high, or at spline junctions where derivatives lose smoothness) the accuracy of the numerical solution is maintained, while allowing a much larger time step to be taken in “easier” sections. Furthermore, Matlab’s ODE solvers have built-in event detection that allows the integration to be terminated precisely at the end of the course. This is an advantage over models such as Carlsson et al. [5], who employed a constant-step Runge-Kutta solver that forced them to choose an unnecessarily small time step so that they could terminate their code as close as possible to the finish line.

### v2

@@ Abstract
Nordic skiing provides fascinating opportunities for mathematical modelling studies that exploit methods and insights from physics, applied mathematics, data analysis, scientific computing and sports science. A typical ski course winds over varied terrain with frequent changes in elevation and direction, and so its geometry is naturally described by a three-dimensional space curve. The skier travels along a course under the influence of various forces, and their dynamics can be described using a nonlinear system of ordinary differential equations (ODEs) that are derived from Newton’s laws of motion. We develop an algorithm for solving the governing equations that combines Hermite spline interpolation, numerical quadrature and a high-order ODE solver. Numerical simulations are compared with measurements of skiers on actual courses to demonstrate the effectiveness of the model. Throughout, we aim to illustrate how elementary concepts from undergraduate courses in calculus and scientific computing can be applied to study real problems in sport, which we hope will provide stimulating examples for both instructors and students. At the same time, we demonstrate how these concepts are capable of providing novel insights into skiing that should also be of interest to sport scientists.
@@ 1 Background
  • We replace the piecewise linear course geometry from MSH [24] with a smoother cubic spline interpolant that more accurately captures the shape of a real ski course (see Fig. 2 for an example).
@@
  • The governing equations are solved using a state-of-the art numerical ODE solver (MATLAB’s ode113) that incorporates high-order derivative approximations, adaptive time stepping, error control and event detection.
@@
  • We incorporate a realistic 3D course (such as in Fig. 1a) that includes track curvature, which allows us to simulate different braking techniques used by athletes to safely negotiate tightly-curved downhill sections [3]. This is important in race settings where athletes strive for any possible advantage over their competitors.
@@
1.1 Exercises, MATLAB Codes and Data Files
@@
All data files and MATLAB codes used to generate the numerical results are also provided with the Supplementary Materials. Note that these codes make use of two functions that are not part of the standard MATLAB distribution: fnder in the Curve Fitting Toolbox [17] and gpxread in the Mapping Toolbox [18].
@@ 2.1 Parameterizing a Ski Course
The path that a skier follows while propelling themselves along a real ski course is a 3D curve such as that depicted in Fig. 1a. This particular plot is generated from GPS data [8] measured on the 4.2 km “Ole” course in Toblach, Italy. This course has been officially certified for competition purposes by FIS (Fédération Internationale de Ski) and has been the location of many high-level races including the 2024 FIS World Cup Tour de Ski. The raw GPS data consists of 58 measurements of latitude $\lambda$, longitude $\psi$, and elevation $z$, which are read using the MATLAB function gpxread and must first be converted to Cartesian coordinates on a flat patch of the Earth’s surface. The curve in Fig. 1a is plotted in terms of $(x,y,z)$, where $z$ is elevation, and $x,y$ lie in the horizontal plane [^5] When latitude $\lambda$ and longitude $\psi$ are measured in degrees, the north-south distance (in m) is $y\approx 111133\psi$ whereas the east-west distance is $x\approx 111413\lambda\cos\left(\pi\lambda/180\right)$. The latter contains an extra cosine adjustment factor because the distance between lines of longitude varies with $\lambda$. These two formulas are actually the leading order terms in an approximation that accounts for the oblate spheroidal shape of the Earth, as explained in the Wikipedia article “Geographic coordinate system.”. For convenience, we also translate coordinates so that the course starts at the origin (highlighted in the plot by a green square). A second view of the course is obtained in Fig. 1b by “unwinding” or straightening this 3D curve and displaying it as a 2D plot of elevation versus a new distance variable $\xi$, which represents the arc length measured along the projected path in the $x,y$-plane (the dotted magenta curve in Fig. 1a). This 2D elevation plot of $(\xi,z)$ is the official method for reporting course elevation profiles in competition documents (private communication, Al Maddox, Canadian FIS Technical Delegate, June 9, 2024).
@@ 2.2 Interpolating with Cubic Hermite Splines
With a bit of foresight, it should be clear that any model of Nordic skiing will require quantities such as slope (or inclination angle) and curvature at the skier’s position $\vec{r}(\alpha)$, parameterized by $\alpha\in[0,1]$. Slope and curvature depend on the first two derivatives of position, $\vec{r}_{\alpha}$ and $\vec{r}_{\alpha\alpha}$, so that any interpolant we use must be twice differentiable. This suggests using a spline interpolant, which is a piecewise-defined function that interpolates each successive pair of data points with a polynomial (refer to the classic text by de Boor for more information [6]). For a 2D course a single interpolant $z=Z(\alpha)$ will suffice, but in 3D we will need to construct three spline interpolants, one for each of the coordinate functions: $x=X(\alpha)$, $y=Y(\alpha)$ and $z=Z(\alpha)$.
@@
In practice, we do not observe any evidence of significant noise or random variation in the course data, and except for certain isolated anomalies discussed later in Section 2.4, the given point locations seem to be generally reliable. This suggests that interpolating splines are a better choice than alternative curve fitting options such as B-splines or other smoothing splines. Nonetheless, we encourage the interested reader to explore the use of other functions like bspline and csaps in MATLAB’s Curve Fitting Toolbox.
@@ 2.3 Computing Distance and Inclination Angle
$\xi(\alpha)=\int_{0}^{\alpha}\sqrt{X_{\alpha}^{2}+Y_{\alpha}^{2}}\,d\alpha$ (1)
@@
$s(\alpha)=\begin{cases}\displaystyle\int_{0}^{\alpha}\sqrt{\Xi_{\alpha}^{2}+Z_{\alpha}^{2}}\;d\alpha,&\text{in 2D,}\\[11.38092pt] \displaystyle\int_{0}^{\alpha}\sqrt{X_{\alpha}^{2}+Y_{\alpha}^{2}+Z_{\alpha}^{2}}\;d\alpha,&\text{in 3D.}\end{cases}$ (2)
To be consistent with other quantities used to represent course geometry, $s$ is also interpolated with a spline to obtain $S(\alpha)$ [^7] If skier position is parameterized by time as $\vec{r}(t)$, then arc length may be expressed elegantly as $s(t)=\int_{0}^{t}\left\|\vec{r}^{\,\prime}(t)\right\|dt=\int_{0}^{t}|v(t)|\,dt$, where $v(t)$ is the skier speed. However, we will not use this formula in our MATLAB implementation..
When implementing these distance calculations in MATLAB, the integrals in (1) and (2) must be computed numerically. To obtain accurate approximations for $s$ (and also $\xi$ for 3D courses), the coordinate splines ($\Xi,Z$ in 2D, $X,Y,Z$ in 3D) are first evaluated on a fine grid of $\alpha$ points. For the 4.2 km Ole course, we use 840 points which yields an average resolution of 5 m for each $\alpha$-interval. The integrals are then approximated using Simpson’s rule, which is fourth-order accurate and has a particularly simple implementation when the $\alpha$ points are equally spaced (see the MATLAB code cumsimpson.m provided with the Supplementary Materials [^8] Conveniently, the spline data structure used in MATLAB permits spline derivatives $X_{\alpha}$, $Y_{\alpha}$, $Z_{\alpha}$ appearing in the arc length integrands to be computed exactly with the function fnder. We initially attempted to use a finite difference approximation for these derivatives, but this generated errors that interfered with the convergence of the ODE solver.). Simpson’s rule is preferable to MATLAB’s built-in trapezoidal rule function trapz because of its much higher accuracy. Once the corresponding fine-grid values of $\xi$ are computed, we can then re-parameterize the curve in terms of $\xi$ by computing a new spline for $Z(\xi)$ (plus $X(\xi)$ and $Y(\xi)$ in 3D). Indeed, the smooth elevation curve plotted in Fig. 1b was generated by first approximating the integral (1) with Simpson’s rule, and then fitting a Hermite spline to the resulting $\xi,z$ data.
@@
Because we have already built the spline $Z(\xi)$ for elevation, the derivative $Z_{\xi}$ can be computed exactly using the fnder function from MATLAB’s Curve Fitting Toolbox. Then $\theta$ can be evaluated at spline points and used to construct another spline $\Theta(\xi)$ for the inclination angle. We have implemented the procedure described above in the MATLAB code setup2d.m, which can be summarized in four main steps:
  • Read the course $\xi,z$ coordinates from the CSV file.
  • Build the spline interpolant $Z(\xi)$ (the default spline type is ’makima’ but can be modified).
  • Compute the spline derivative $Z_{\xi}$ and use it to construct splines $S(\xi)$ and $\Theta(\xi)$.
  • Print diagnostics regarding point spacing, arclength and course gradient for homologation purposes.
More details are provided in Section 5 regarding the corresponding code setup3d.m for constructing a 3D parametric spline representation from GPS data.
@@
The model devloped by MSH [24] parameterizes a ski course as $z(s)$ using arc length $s$, and replaces slope $\frac{dz}{d\xi}$ in Eq. (3) with $\frac{dz}{ds}\big/\sqrt{1-(\frac{dz}{ds})^{2}}$. Show that these two expressions for slope are equivalent for a “reasonable” ski course with no self-intersections and finite slope.
@@ 2.4 Homologation and Data Errors
All ski courses that run high-level Nordic races must adhere to guidelines set out in the FIS Cross-Country Homologation Manual [10]. In the context of skiing, homologation refers to the FIS-certified process that any race course must go through before an official competition can be held. The Manual sets out constraints on course terrain and stadium layout, including limits on the number, length, and slope of hills that are allowed [^9] An easily-digestible summary of the homologation rules can be found on the Wikipedia page “Cross-country skiing trail”.. These specifications are implemented in the form of an Excel spreadsheet that FIS officials use to perform the necessary calculations [10].
@@
    To ensure athlete safety, a maximum limit on curvature of downhill sections is imposed in terms of centripetal acceleration, $a_{c}=v^{2}/R$, where $v$ is the skier speed and $R$ is the radius of curvature. A table in [10, p. 26] identifies various limits on $a_{c}$ ranging from 2–25 m/s2 depending on race format.
@@ 3.1 ODE System for Skier Dynamics
    where $m$ is the skier mass (in kg) and $P(v)$ represents the power (in watts or kg m2/s3) that is exerted by the muscles to generate forward motion. The $\frac{1}{v}$” dependence implies that the skier works hardest to accelerate at low speeds, whereas the power demands on their muscles ease off at higher speeds [^11] Our earlier assumption that $v>0$ ensures that this propulsion force is well-defined. However, we could handle a zero speed by replacing the factor $v$ in the denominator of (4) with $\max(v,\varepsilon)$ for some small cut-off parameter $\varepsilon>0$.. This function was one of three different functional forms proposed by MSH [24] and fit to power measurements of Nordic ski racers, and the specific form for $P(v)$ will be described later in Section 3.3.
@@
  • Aerodynamic drag force $F_{d}(v)=-\beta v^{2}$: which we assume depends quadratically on the speed relative to still air [^12] If there is a significant headwind/tailwind, then $v$ should be replaced with a wind-corrected speed (see Exercise 16).. The parameter $\beta=\frac{\rho C_{d}A}{2m}$ (units of m-1) is proportional to air density $\rho$ (in kg/m3), skier cross-sectional area $A$ (m2), and a dimensionless drag coefficient $C_{d}$. As their speed varies throughout a race, trained skiers will adjust their technique and body orientation so that the area $A$ varies with the skier’s speed. So in practice, we will consider the product $C_{d}A(v)$ to be a given function of speed (discussed in more detail in Section 3.3).
@@
The inclination angle appearing in the right-hand side of both ODEs can be obtained by simply evaluating the spline $\theta=\Theta(\xi)$ constructed in Section 2.3.
@@
Use dimensional analysis (by applying the Buckingham Pi Theorem [29]) to derive a general form for the propulsion force in Eq. (4), assuming only that $F_{p}$ depends on the quantities ${P}_{\text{\emph{max}}}$, $m$ and $v$.
@@
    Find an explicit formula for speed $v(t)$ when the course profile is $z(\xi)=b\xi$ for some constant slope $b$. Assume that the skier applies a constant propulsion force, $F_{p}(v)\equiv F_{po}$, which implies that the MSH power function is linear in the speed, $P(v)\propto v$. In constrast with Eq. (4), this force has no $\frac{1}{v}$” singularity, so you may also assume that the skier starts from rest with $v(0)=0$.
@@
    Plot the speed $v(t)$ for an uphill climb with a gradient of ++10% ($b=0.1$) and assume parameter values $\beta=0.0045$, $g=9.8$, $\mu=0.037$, and $F_{po}=1.4$ from the MSH baseline test (see Table 1 and Fig. 5). Observe that the skier’s speed initially increases but then gradually flattens out over time and approaches a constant terminal value. Estimate this terminal speed and the time required to reach it.
@@
    Repeat part (b) on a constant downhill gradient of -10%, assuming the skier rests in a tuck position with $\beta=0.0025$ and $F_{po}=0$. Compare your results for the uphill and downhill slopes.
@@ 3.3 Fitting Parameters to Skier Measurements
The final remaining detail is the locomotive power function $P(v)$ introduced in Eq. (4). MSH used extensive measurements of ski racers to propose three different functional forms for $P(v)$, which they compared using numerical simulations. We will adopt an exponential function which they refer to as “Model 1” and used for their baseline simulations [^14] The power functions corresponding to MSH Models 2,3 have a similar shape to the baseline, and don’t have a major impact on the resulting solution, something the interested reader is welcome to explore.:
$P(v)={P}_{\text{\emph{max}}}\exp\big(\text{--}\,(v/b)^{n}\big).$ (10)
This function is designed to mimic an athlete who generates the highest power when skiing uphill (at lower speeds) and the lowest on downhills (at high speeds). This reflects the same adjustments in technique we discussed for the drag coefficient, where the skier has to use their entire body to work hardest on steep uphill sections and maintain their speed, whereas on steep downhills they shift into a tuck position to conserve energy and minimize drag. An intermediate level of power is expended on undulating terrain, where a relatively high speed can be maintained using less energy-intensive techniques. MSH estimated the parameters in this power equation by fitting with experimental data to obtain $n=4$, $b=10$, and ${P}_{\text{\emph{max}}}=275$. These authors indicate that their maximum power value was obtained from measurements of an elite athlete who was “skiing at a self-chosen pace.” We will see later in comparison with another set of experiments by Welde on elite racers (Section 4.2) that even though this athlete’s pacing might be considered fast, they were almost surely not skiing at a race pace consistent with elite skiers.
The next section presents several examples of simulations that start from skiers with the baseline parameter values, and explore how changes in certain parameters impact the solution. A more comprehensive parameter sensitivity study can be found in MSH.
@@
The system of nonlinear ODEs (5)–(6), or its augmented formulation (5)–(8), can be solved easily and accurately using MATLAB’s variable-step, variable- order ODE solver ode113. This solver employs Adams-Bashforth-Moulton formulas with derivative approximations having orders from 1 to 13, adaptively selecting both the order and the time step so as to minimize computational cost while also satisfying the given error tolerances. A comprehensive comparison of this and other MATLAB ODE solvers can be found in the paper [30], and Exercise 12 provides a chance to compare the performance of several different solvers.
This approach has been implemented in our code skirun2d.m with the parameter neqs (the number of ODEs) set to either 2 or 4 to select the desired problem formulation. In the following sections, we apply the code to solving three test examples:
@@
    The MSH baseline simulation [24] on a 4.2 km course using the reduced (2-ODE) model. This provides a convenient test to validate our MATLAB code and investigate any differences that arise due to our choice of ODE solver and spline interpolant (linear or cubic Hermite).
@@
A major advantage of ode113 is that it adaptively adjusts the time step and the order of the numerical method to satisfy user-specified error tolerances (we use AbsTol = RelTol = 1e-8). This ensures that especially in the more difficult sections of the course (where curvature is high, or at spline junctions where derivatives lose smoothness) the accuracy of the numerical solution is maintained, while allowing a much larger time step to be taken in “easier” sections. Furthermore, MATLAB’s ODE solvers have built-in event detection that allows the integration to be terminated precisely at the end of the course. This is an advantage over models such as Carlsson et al. [5], who employed a constant-step Runge-Kutta solver, which forced them to choose an unnecessarily small time step so that they could terminate their code as close as possible to the finish line.